40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 15 03:08:57 2015, ericq, Update, LSC, DRFPMI Locked for 20 sec DRFPMI_locked.pdf
    Reply  Thu Oct 15 04:14:14 2015, ericq, Update, LSC, DRFPMI Locked for 20 sec 
    Reply  Thu Oct 15 10:59:12 2015, Koji, Update, LSC, DRFPMI Locked for 20 sec 
    Reply  Sat Oct 17 18:55:07 2015, rana, Update, LSC, DRFPMI Locked for 20 sec 
       Reply  Mon Oct 19 15:23:22 2015, ericq, Update, LSC, Longer DRFPMI lock second_DRFPMI_15Oct2015.png
          Reply  Tue Oct 20 11:24:29 2015, ericq, HowTo, LSC, How to DRFPMI 
Message ID: 11698     Entry time: Mon Oct 19 15:23:22 2015     In reply to: 11696     Reply to this: 11701
Author: ericq 
Type: Update 
Category: LSC 
Subject: Longer DRFPMI lock 

Here is a longer lock, about 100 seconds RF only, from later that same night. The in-loop CARM and DARM error signals have the order of magnitude of 1nm per count. 

From ~-150 to -103, we were fine tuning the ALS offsets to try and get close to the real CARM/DARM zero points then blending the RF CARM signal. 

At -100, the CARM bandwidth increases to a few kHz and stabilizes the arm powers. By -81, the error signals are all RF. At -70, I turned on the transmon QPD servos, which brought the power up a bit. 

If I recall correctly, lock was lost because I put waaaay too big of an excitation on DARM with the goal of running its UGF servo for a bit. The number I entered was appropriate for ALS, but most certainly too huge for AS55...

ELOG V3.1.3-